User Tools

Site Tools


handover:r4711

Differences

This shows you the differences between two versions of the page.

Link to this comparison view

Next revision
Previous revision
handover:r4711 [2015/10/29 18:32]
Ross Turner
handover:r4711 [2015/10/30 16:00]
Ross Turner
Line 3: Line 3:
 **Hobart 26m** **Hobart 26m**
  
-Module: HAY-005, start pos 1895.329 GB+Module: HAY-005, start pos 1895.329 GB
  
-     * 18:15 UT - wth command not working, gives error "​Problem connecting to metserver socket. Illegal error class code: 0xffffffff."​ I had a look at the station file in pfmed but it looked. Not urgent so I'll report it at a more reasonable hour. (Ellen)+     * 18:15 UT - wth command not working, gives error "​Problem connecting to metserver socket. Illegal error class code: 0xffffffff."​ I had a look at the station file in pfmed but nothing I tried worked. Not urgent so I'll report it at a more reasonable hour. (Ellen)
      * 18:31 UT - Experiment started ok. (Ellen)      * 18:31 UT - Experiment started ok. (Ellen)
 +     * 19:49 UT - Again, I am getting alarms fairly regularly, saying "ERROR m5 -104 mk5cn: time-out, connection closed"​. This doesn'​t appear to mean anything as everything still seems fine to me. It has also happened at 19:51, 20:19, 20:22, 21:10, 21:13 UT. (Ellen)
 +     * 22:15 UT - Had to halt schedle for a quick maintinence check. ​ Also, had to restart the field system to reset some settings. ​ The timeout on the Mark 5A was set to 1/10 of a second, which was occasionally setting off a time out alarm. ​ This problem has now been fixed. ​ Schedule restarted and recording to a new module. ​ Disk pos cannot be checked due to change of module. (Imogen)
 +     * The weather monitor machine has failed and is being fixed today. ​ Weather is being reported manually via the BOM website. (Imogen)
 +     * 04:02UT Schedule halted due to reciever maintence (Imogen)
 +     * 04:34UT schedule resumed after reciever maintence (Imogen)
 +
 +     * 0817 UT Telescope started recording to HOB+0095 at some time during Imogen'​s shift; this wasn't noted. Currently at 399 GB on this module (JS)
 +     * 0822 UT 400.8 GB recorded to HOB+0095 of 650 GB total for experiment to this point. I'm guessing HAY-005 was a ~2TB module and a deficit accounts for the rest (JS)
 +     * 0824 UT There is ~ 150 GB of missing data (I think, i don't know when Imogen changed modules) but the scans that were missed were not noted (sorry Ross) (JS)
  
 **Katherine** **Katherine**
Line 13: Line 22:
  
      * 18:30 UT - Experiment started ok. (Ellen)      * 18:30 UT - Experiment started ok. (Ellen)
 +     * 15:54 UT - Antenna stuck, scans 303-1553 and 303-1558 unreliable (Ross).
  
  
/home/www/auscope/opswiki/data/pages/handover/r4711.txt · Last modified: 2015/10/30 16:00 by Ross Turner